# A survey of degree-boundedness

### v1

@@ 3.2 Minimum degree and clique number
There are also many nice geometric examples of graph classes that are degree-bounded by $\omega$. For each integer $d$, the class of all intersection graphs of balls in $\mathbb{R}^{d}$ is degree-bounded by $\omega$. (The intersection graph of a finite collection $\mathcal{B}$ of balls in $\mathbb{R}^{d}$ is the graph with vertex set $\mathcal{B}$, where two vertices are adjacent if they have non-empty intersection as balls in $\mathbb{R}^{d}$.) In fact, this holds if instead of balls we consider compact convex sets in $\mathbb{R}^{d}$ with bounded aspect ratio, following from work on separators [74, 100, 122]. Dvořák, Norin, and the second author [47] also showed that we can allow the aspect ratio to be unbounded as long as for any pair of shapes $A,B\subseteq\mathbb{R}^{d}$ under consideration, we can “smoothly slide one of $A,B$ around the interior of the other”. For instance, we can allow $A$ and $B$ to be axis-aligned rectangles as long as $B$ contains some translate of $A$ or vice-versa. We note that for all of these examples, it can actually be shown that the shape of smallest volume has small minimum degree (as a function of $\omega$).
By combining Theorem 1.1 of Girão and Hunter [67] with well-known bounds on off-diagonal Ramsey numbers [2], we obtain the following result. Note that by considering a greedy coloring, Corollary 3.5 implies that every hereditary class which is degree-bounded by $\omega$ is also polynomially $\chi$-bounded. This is a special case of Esperet’s Conjecture.
@@
3.3 The Erdös-Hajnal Conjecture
We already discussed the connection between Esperet’s Conjecture and the Erdös-Hajnal Conjecture in the last subsection. Now let us discuss the connection between the Erdös-Hajnal Conjecture and Theorem 1.1 about degree-boundedness.
@@ 3.3 The Erdös-Hajnal Conjecture
Let us note right away that Conjecture 3.6 is actually a special case of the Erdös-Hajnal Conjecture and so was already conjectured in [50]. In fact, we now use Theorem 1.1 to argue that Conjecture 3.6 is equivalent to the Erdös-Hajnal Conjecture for hereditary, degree-bounded classes.
@@
A hereditary, degree-bounded class $\mathcal{F}$ has the Erdös-Hajnal property if and only if there exists a polynomial $\widehat{p}$ so that every graph $G\in\mathcal{F}$ satisfies $\delta(G)\leqslant\widehat{p}(\widehat{\tau}(G))$.
@@
Note that, in order to prove that $\mathcal{F}$ has the Erdös-Hajnal property, it is enough to prove that every $n$-vertex graph in $\mathcal{F}$ has a clique or an independent set of size $\Omega(n^{\epsilon})$. So we may assume that $n$ is sufficiently large, and thus
@@
So indeed, the class $\mathcal{F}$ has the Erdös-Hajnal property.
In the other direction, suppose that $\mathcal{F}$ is a hereditary, degree-bounded class which has the Erdös-Hajnal property. Let $\epsilon>0$ be such that every $n$-vertex graph $G\in\mathcal{F}$ has a clique or an independent set of size at least $n^{\epsilon}$. Then every graph $G\in\mathcal{F}$ with biclique number $\tau=\tau(G)$ has induced biclique number at least $\tau^{\epsilon}$, for we can transform each side of the biclique into either a clique or an independent set of size $\tau^{\epsilon}$. Thus $\widehat{\tau}(G)\geqslant\tau(G)^{\epsilon}$ for every $G\in\mathcal{F}$.
@@
We recall from the earlier discussion that Nguyen, Scott, and Seymour [105] recently proved that every class of bounded VC-dimension has the Erdös-Hajnal Property. This implies Conjecture 3.6 for many interesting cases. They also proved in [103] that the class of graphs without an induced subgraph isomorphic to the $5$-vertex path has the Erdös-Hajnal Property. This theorem resolved a case that had been open for a long time and received a lot of attention. However, it is still open whether the class of graphs with no induced subgraph isomorphic to the $6$-vertex path has the Erdös-Hajnal Property, and this class is degree-bounded by [80]. So Conjecture 3.6 seems quite difficult.
@@ 3.4 Direct implications
We already argued that this class is $\chi$-bounded. To see that it is not degree-bounded, we combine two theorems. First, in order to disprove another conjecture about $\chi$-boundedness, Pawlik, Kozik, Krawczyk, Lasoń, Micek, Trotter, and Walczak [107] proved that Burling graphs can be represented by the intersection pattern of a collection of line segments in the plane. That is, they proved that for any Burling graph $G_{k}$, there is a collection $\{L_{v}:v\in V(G_{k})\}$ of line-segments in the real plane $\mathbb{R}^{2}$ so that for any pair of distinct vertices $u,v\in V(G_{k})$, the vertices $u$ and $v$ are adjacent in $G_{k}$ if and only if the line segments $L_{v}$ and $L_{u}$ intersect. Second, Fox and Pach [54] proved that the class of all graphs with such a representation is degree-bounded. (In fact, they proved this even for the more general class of string graphs which will be discussed in Subsection 5.1.) Since the class of all graphs with such a representation is hereditary, Theorem 3.8 follows.
@@
We write $o_{\mathcal{F}}(1)$ for a function whose limit as $n$ goes to infinity is 00, when the class $\mathcal{F}$ is fixed.
The intersection graphs mentioned before – string graphs – satisfy an even stronger bound. Generalizing work of McGuinness [99], Fox and Pach [56] showed that every $n$-vertex, triangle-free string graph has chromatic number at most polylogarithmic in $n$. We note that there is a long history of studying the maximum chromatic number of $n$-vertex, triangle-free graphs with geometric representations; see also [58, 124, 132] for other results.
@@ 3.5 Perfect graphs
A graph $G$ is perfect if for every induced subgraph $H$ of $G$, the chromatic number of $H$ is equal to its clique number, that is, $\chi(H)=\omega(H)$. Berge [8] conjectured a wonderful characterization of perfect graphs: that a graph is perfect if and only if it does not have an induced subgraph which is a cycle of length odd and greater than $3$, or the complement of such a graph. (The length of a cycle is the number of edges it contains.) Chudnovsky, Robertson, Seymour, and Thomas [26] proved this conjecture in 2006, for which they won the famous Fulkerson Prize in discrete mathematics. Their theorem is known as the Strong Perfect Graph Theorem; the Weak Perfect Graph Theorem is due to Lovász [93] and says that a graph is perfect if and only if its complement is perfect. Perfect graphs also have strong connections to combinatorial optimization. They can be used to characterize when the polytope of a set packing problem has only integral vertices; see the book by Cornuéjols [31].
In our setting, it is natural to propose the following analog of perfect graphs. We say that a graph $G$ is degree-perfect if for every induced subgraph $H$ of $G$, the minimum degree of $H$ is equal to its biclique number, that is, $\delta(H)=\tau(H)$. Together with the standard characterization of bipartite graphs, a theorem of Golumbic and Goss [68, Theorem 4] implies the following characterization of degree-perfect graphs.
@@
We note that the class of graphs considered in Theorem 3.11 is of independent interest; they are the “bipartite analog” of chordal graphs. Recall that a graph is chordal if its only induced cycles are copies of the triangle $C_{3}$. Theorem 3.11 equivalently says that degree-perfect graphs are precisely the bipartite graphs whose only induced cycles are copies of the $4$-cycle $C_{4}$. The theorem of Golumbic and Goss [68, Theorem 4] which implies Theorem 3.11 says that such graphs have simplicial edges. Simplicial edges are defined analogously to the simplicial vertices which are studied in chordal graphs; an edge $e=\{u,v\}$ of a bipartite graph $G$ is simplicial if $G$ contains all possible edges between $N(u)$ and $N(v)$. See [68] for details.
@@ 3.6 Induced subdivisions
It was originally conjectured that for every graph $H$, the class $\mathcal{F}_{H}$ is $\chi$-bounded [114]. This problem has some nice motivation. First of all, Scott [114] proved that it is true when $H$ is a tree (that is, that the class $\mathcal{F}_{T}$ is $\chi$-bounded for any tree $T$). Moreover, one of Gyárfás’ [72] original problems on $\chi$-boundedness from the 80’s equivalently says that the class $\mathcal{F}_{C}$ is $\chi$-bounded for any cycle $C$. This conjecture was later proven by Chudnovsky, Scott, and Seymour [27]. (In fact, with Spirkl [29], they also showed that it is enough to forbid odd induced subdivisions of a cycle, that is, where each path $P_{e}$ used to replace an edge $e$ of the cycle has odd length.)
@@
We would also like to point the reader towards the following conjecture of Scott and Seymour [116], which would combine some of the topics discussed here.
@@
A subdivision of $H$ is odd if for each edge $e\in E(H)$, the path $P_{e}$ used to replace $e$ has odd length. We refer the reader to [116] for a discussion on the progress towards Conjecture 3.13.
@@ 3.7 Induced trees
The most famous problem on $\chi$-boundedness which is still open is the Gyárfás-Sumner Conjecture [70, 125], which proposes that for any forest $T$, the class of all graphs $G$ with no induced subgraph isomorphic to $T$ is $\chi$-bounded. (It is known that forests are the only graphs that can have this property, as there exist graphs of arbitrarily large girth and chromatic number; see for instance the classical random construction of Erdős [49].) There is a lot of partial progress on this conjecture; see [116]. Notably, Gyárfás [72, Theorem 2.4] proved the conjecture for paths using a now famous path-augmentation argument.
While the Gyárfás-Sumner Conjecture remains open, its degree-bounded analog has been known for a long time, as discussed in the survey of Scott and Seymour [116]. First of all, Hajnal and Rödl independently proved that for any forest $T$ and any integer $t$, the class of all graphs $G$ with no induced subgraph isomorphic to $T$ and with $\tau(G)\leqslant t$ is $\chi$-bounded; see [73]. Kierstead and Penrice [80, Corollary 7] then proved a theorem which implies the following.
Theorem 3.14.
[80] For any forest $T$, the class of all graphs with no induced subgraph isomorphic to $T$ is degree-bounded.
@@ 4 Degree-bounding functions
In this section, we discuss the rate of growth of optimal degree bounding functions. As mentioned before, in contrast to $\chi$-boundedness, every hereditary degree-bounded class $\mathcal{F}$ has a degree-bounding function that is a polynomial. In this section we give a broad overview of how to prove this sort of theorem.
The key idea is that the minimum degree of the $4$-cycle free graphs in the class “controls” the rate of growth of the whole function. Throughout this section, we let $\mathcal{F}$ denote an arbitrary hereditary degree-bounded class. Recall from Lemma 2.1 that bounding the minimum degree is the same thing as bounding the average degree. Thus, there exists an integer $k:=k({\mathcal{F}})$ such that every $4$-cycle-free graph in $\mathcal{F}$ has average degree less than $k$. The authors, together with Girão, Hunter, and Scott [44], proved that every hereditary degree-bounded class has a degree bounding function that is singly exponential in the biclique number $\tau$.
Theorem 4.1.
[44] There exists a universal constant $C$ so that $f_{k}(\tau)=k^{C\tau^{3}}$ is a degree-bounding function for any hereditary degree-bounded class $\mathcal{F}$.
@@
There exists a universal constant $C$ so that $p_{k}(s)=\tau^{Ck^{4}}$ is a degree-bounding function for any hereditary degree-bounded class $\mathcal{F}$.

### v2

@@ 3.2 Minimum degree and clique number
There are also many nice geometric examples of graph classes that are degree-bounded by $\omega$. For each integer $d$, the class of all intersection graphs of balls in $\mathbb{R}^{d}$ is degree-bounded by $\omega$. (The intersection graph of a finite collection $\mathcal{B}$ of balls in $\mathbb{R}^{d}$ is the graph with vertex set $\mathcal{B}$, where two vertices are adjacent if they have non-empty intersection as balls in $\mathbb{R}^{d}$.) In fact, this holds if instead of balls we consider compact convex sets in $\mathbb{R}^{d}$ with bounded aspect ratio, following from work on separators [79, 106, 131]. Dvořák, Norin, and the second author [49] also showed that we can allow the aspect ratio to be unbounded as long as for any pair of shapes $A,B\subseteq\mathbb{R}^{d}$ under consideration, we can “smoothly slide one of $A,B$ around the interior of the other”. For instance, we can allow $A$ and $B$ to be axis-aligned rectangles as long as $B$ contains some translate of $A$ or vice-versa. We note that for all of these examples, it can actually be shown that the shape of smallest volume has small degree (as a function of $\omega$).
By combining Theorem 1.1 of Girão and Hunter [72] with well-known bounds on off-diagonal Ramsey numbers [3], we obtain the following result. Note that by considering a greedy/degeneracy coloring, Corollary 3.5 implies that every hereditary class which is degree-bounded by $\omega$ is also polynomially $\chi$-bounded. This is a special case of Esperet’s Conjecture.
@@
3.3 The Erdős-Hajnal Conjecture
We already discussed the connection between Esperet’s Conjecture and the Erdős-Hajnal Conjecture in the last subsection. Now let us discuss the connection between the Erdős-Hajnal Conjecture and Theorem 1.1 about degree-boundedness.
@@ 3.3 The Erdős-Hajnal Conjecture
Let us note right away that Conjecture 3.6 is actually a special case of the Erdős-Hajnal Conjecture and so was already conjectured in [52]. In fact, we now use Theorem 1.1 to argue that Conjecture 3.6 is equivalent to the Erdős-Hajnal Conjecture for hereditary, degree-bounded classes.
@@
A hereditary, degree-bounded class $\mathcal{F}$ has the Erdős-Hajnal property if and only if there exists a polynomial $\widehat{p}$ so that every graph $G\in\mathcal{F}$ satisfies $\delta(G)\leqslant\widehat{p}(\widehat{\tau}(G))$.
@@
Note that, in order to prove that $\mathcal{F}$ has the Erdős-Hajnal property, it is enough to prove that every $n$-vertex graph in $\mathcal{F}$ has a clique or an independent set of size $\Omega(n^{\epsilon})$. So we may assume that $n$ is sufficiently large, and thus
@@
So indeed, the class $\mathcal{F}$ has the Erdős-Hajnal property.
In the other direction, suppose that $\mathcal{F}$ is a hereditary, degree-bounded class which has the Erdős-Hajnal property. Let $\epsilon>0$ be such that every $n$-vertex graph $G\in\mathcal{F}$ has a clique or an independent set of size at least $n^{\epsilon}$. Then every graph $G\in\mathcal{F}$ with biclique number $\tau=\tau(G)$ has induced biclique number at least $\tau^{\epsilon}$, for we can transform each side of the biclique into either a clique or an independent set of size $\tau^{\epsilon}$. Thus $\widehat{\tau}(G)\geqslant\tau(G)^{\epsilon}$ for every $G\in\mathcal{F}$.
@@
We recall from the earlier discussion that Nguyen, Scott, and Seymour [112] recently proved that every class of bounded VC-dimension has the Erdős-Hajnal Property. This implies Conjecture 3.6 for many interesting cases. They also proved in [110] that the class of graphs without an induced subgraph isomorphic to the $5$-vertex path has the Erdős-Hajnal Property. This theorem resolved a case that had been open for a long time and received a lot of attention. However, it is still open whether the class of graphs with no induced subgraph isomorphic to the $6$-vertex path has the Erdős-Hajnal Property, and this class is degree-bounded by [86]. So Conjecture 3.6 seems quite difficult. We note that Nguyen, Scott, and Seymour [109] recently proved that classes with a forbidden induced path “almost” have the Erdős-Hajnal Property; they have cliques or independent sets of size $2^{(\log n)^{1-o(1)}}$.
@@ 3.4 Direct implications
We already argued that this class is not $\chi$-bounded. To see that it is degree-bounded, we combine two theorems. First, in order to disprove another conjecture about $\chi$-boundedness, Pawlik, Kozik, Krawczyk, Lasoń, Micek, Trotter, and Walczak [116] proved that Burling graphs can be represented by the intersection pattern of a collection of line segments in the plane. That is, they proved that for any Burling graph $G_{k}$, there is a collection $\{L_{v}:v\in V(G_{k})\}$ of line-segments in the real plane $\mathbb{R}^{2}$ so that for any pair of distinct vertices $u,v\in V(G_{k})$, the vertices $u$ and $v$ are adjacent in $G_{k}$ if and only if the line segments $L_{v}$ and $L_{u}$ intersect. Second, Fox and Pach [57] proved that the class of all graphs with such a representation is degree-bounded. (In fact, they proved this even for the more general class of string graphs which will be discussed in Subsection 5.1.) Since the class of all graphs with such a representation is hereditary, Theorem 3.8 follows.
@@
We write $o_{\mathcal{F}}(1)$ for a function whose limit as $n$ goes to infinity is 00, when the class $\mathcal{F}$ is fixed. We note that Problem 3.9 equivalently asks if every $n$-vertex, triangle-free graph in $\mathcal{F}$ has an independent set of size at least $n^{1-o_{\mathcal{F}}(1)}$. (Since the class $\mathcal{F}$ is hereditary, the desired coloring can then be obtained by greedily removing a maximum independent set.)
In upcoming work, Nguyen, Scott, and Seymour [114] prove that for each fixed tree $H$, every $n$-vertex, triangle-free graph with no induced subgraph isomorphic to $H$ has chromatic number at most $n^{o_{H}(1)}$. The intersection graphs mentioned before – string graphs – satisfy an even stronger bound. Generalizing work of McGuinness [105], Fox and Pach [59] showed that every $n$-vertex, triangle-free string graph has chromatic number at most polylogarithmic in $n$. In upcoming work, Nguyen, Scott, and Seymour [113] generalize this theorem to graphs with a forbidden induced subdivision of any fixed graph $H$. All of these results also hold for graphs which are $K_{t}$-free, rather than triangle-free. Finally, we note that there is a long history of studying the maximum chromatic number of $n$-vertex, triangle-free graphs with geometric representations; see also [61, 134, 143] for other results.
@@ 3.5 Perfect graphs
A graph $G$ is perfect if for every induced subgraph $H$ of $G$, the chromatic number of $H$ is equal to its clique number, that is, $\chi(H)=\omega(H)$. Berge [10] conjectured a wonderful characterization of perfect graphs: that a graph is perfect if and only if it does not have an induced subgraph which is a cycle of length odd and greater than $3$, or the complement of such a graph. (The length of a cycle or a path is the number of edges it contains.) Chudnovsky, Robertson, Seymour, and Thomas [28] proved this conjecture in 2006, for which they won the famous Fulkerson Prize in discrete mathematics. Their theorem is known as the Strong Perfect Graph Theorem; the Weak Perfect Graph Theorem is due to Lovász [99] and says that a graph is perfect if and only if its complement is perfect. Perfect graphs also have strong connections to combinatorial optimization. They can be used to characterize when the polytope of a set packing problem has only integral vertices; see the book by Cornuéjols [33].
In our setting, it is natural to propose the following analog of perfect graphs. We say that a graph $G$ is degree-perfect if for every induced subgraph $H$ of $G$, the minimum degree of $H$ is at most its biclique number, that is, $\delta(H)\leqslant\tau(H)$. Together with the standard characterization of bipartite graphs, a theorem of Golumbic and Goss [73, Theorem 4] implies the following characterization of degree-perfect graphs.
@@
We note that the class of graphs considered in Theorem 3.11 is of independent interest; they are the “bipartite analog” of chordal graphs. Recall that a graph is chordal if its only induced cycles are copies of the triangle $C_{3}$. Theorem 3.11 equivalently says that degree-perfect graphs are precisely the bipartite graphs whose only induced cycles are copies of the $4$-cycle $C_{4}$. The theorem of Golumbic and Goss [73, Theorem 4] which implies Theorem 3.11 says that such graphs have simplicial edges. Simplicial edges are defined analogously to the simplicial vertices which are studied in chordal graphs; an edge $e=\{u,v\}$ of a bipartite graph $G$ is simplicial if $G$ contains all possible edges between $N(u)$ and $N(v)$. See [73] for details.
@@ 3.6 Induced subdivisions
It was originally conjectured that for every graph $H$, the class $\mathcal{F}_{H}$ is $\chi$-bounded [123]. This problem has some nice motivation. First of all, Scott [123] proved that it is true when $H$ is a tree (that is, that the class $\mathcal{F}_{T}$ is $\chi$-bounded for any tree $T$). Moreover, one of Gyárfás’ [77] original problems on $\chi$-boundedness from the 80’s equivalently says that the class $\mathcal{F}_{C}$ is $\chi$-bounded for any cycle $C$. This conjecture was later proven by Chudnovsky, Scott, and Seymour [29]. (In fact, with Spirkl [31], they also showed that it is enough to forbid all induced odd cycles of length at least $\ell$, for any fixed integer $\ell$. A cycle is odd if it has an odd number of edges.)
@@
We would also like to point the reader towards the following conjecture of Scott and Seymour [125], which combines some of the topics discussed here. We say that a subdivision of a graph $H$ is odd (respectively, even) if for each edge $e\in E(H)$, the path $P_{e}$ used to replace $e$ contains an odd (respectively, even) number of edges.
@@
We refer the reader to [125] for a discussion on the progress towards Conjecture 3.13.
We note that for any graph $H$ which contains an odd cycle, the class of all graphs with no induced odd subdivision of $H$ is not degree bounded, because it contains all bipartite graphs. For even subdivisions, however, we obtain the following result as a corollary of Theorem 3.12 and Theorem 2.6.
Corollary 3.14.
For any graph $H$, the class of all graphs with no induced even subdivision of $H$ is degree-bounded.
Proof.
By Theorem 2.6, it suffices to prove that the class of all bipartite graphs with no induced even subdivision of $H$ is degree-bounded. We claim that this class also excludes all induced subdivisions of a clique on $(2|V(H)|-1)$-many vertices. So, suppose that $G$ is a bipartite graph which has an induced subdivision of a clique on $(2|V(H)|-1)$-many vertices. Note that for any two vertices $u$ and $v$ in the same side of $G$, every path between $u$ and $v$ has an even number of edges. So by the pigeonhole principle, $G$ contains an induced even subdivision of a clique on $|V(H)|$-many vertices, and thus an induced even subdivision of $H$.
The corollary now follows from Theorem 3.12. ∎
@@ 3.7 Induced trees
The most famous problem on $\chi$-boundedness which is still open is the Gyárfás-Sumner Conjecture [75, 135], which proposes that for any forest $T$, the class of all graphs $G$ with no induced subgraph isomorphic to $T$ is $\chi$-bounded. (It is known that forests are the only graphs that can have this property, as there exist graphs of arbitrarily large girth and chromatic number; see for instance the classical random construction of Erdős [51].) There is a lot of partial progress on this conjecture; see [125]. Notably, Gyárfás [77, Theorem 2.4] proved the conjecture for paths using a now famous path-augmentation argument.
While the Gyárfás-Sumner Conjecture remains open, its degree-bounded analog has been known for a long time, as discussed in the survey of Scott and Seymour [125]. First of all, Hajnal and Rödl independently proved that for any forest $T$ and any integer $t$, the class of all graphs $G$ with no induced subgraph isomorphic to $T$ and with $\tau(G)\leqslant t$ is $\chi$-bounded; see [78]. Kierstead and Penrice [86, Corollary 7] then proved a theorem which implies the following.
Theorem 3.15 ([86]).
For any forest $T$, the class of all graphs with no induced subgraph isomorphic to $T$ is degree-bounded.
@@ 4 Degree-bounding functions
In this section, we discuss the rate of growth of optimal degree-bounding functions. As mentioned before, in contrast to $\chi$-boundedness, Girão and Hunter [72] proved that every hereditary degree-bounded class has a degree-bounding function that is a polynomial. In this section we give a broad overview of how to prove this sort of theorem.
The key idea is that the minimum degree of the $4$-cycle free graphs in the class “controls” the rate of growth of the whole function. Throughout this section, we let $\mathcal{F}$ denote an arbitrary hereditary degree-bounded class. Recall from Lemma 2.1 that bounding the minimum degree is the same thing as bounding the average degree. Thus, there exists an integer $k:=k({\mathcal{F}})$ such that every $4$-cycle-free graph in $\mathcal{F}$ has average degree less than $k$. The authors, together with Girão, Hunter, and Scott [46], proved that every hereditary degree-bounded class has a degree-bounding function that is singly exponential in the biclique number $\tau$.
Theorem 4.1 ([46]).
There exists a universal constant $C$ so that $f_{k}(\tau)=k^{C\tau^{3}}$ is a degree-bounding function for any hereditary degree-bounded class $\mathcal{F}$.
@@
There exists a universal constant $C$ so that $p_{k}(\tau)=\tau^{Ck^{4}}$ is a degree-bounding function for any hereditary degree-bounded class $\mathcal{F}$.